OpenStudy (anonymous):

\[\sum _{k=0}^{\infty } (-1)^{k+1}\left(\frac{1}{3}\right)^k\]

OpenStudy (anonymous):

What does it converge to? I know how to do it , want to know other ways

OpenStudy (anonymous):

\[=\sum_{0}^{\infty}(-1)*(-1)^{k}*\left(\frac{1}{3}\right)^{k}\]\[=\sum_{0}^{\infty}(-1)*\left(-\frac{1}{3}\right)^{k}\]\[= -\sum_{0}^{\infty}\left(-\frac{1}{3}\right)^{k}\]\[= -\left(\frac{3}{1+3}\right)\]\[= -\frac{3}{4}\] That's how I'd do it.

OpenStudy (aravindg):

imru i posted queston

OpenStudy (anonymous):

That's how I did it too

OpenStudy (aravindg):

...........